# Cross-Modal Conditioned Reconstruction for Language-guided Medical Image Segmentation

### v1

@@ Abstract
Recent developments underscore the potential of textual information in enhancing learning models for a deeper understanding of medical visual semantics. However, language-guided medical image segmentation still faces a challenging issue. Previous works employ implicit and ambiguous architectures to embed textual information. This leads to segmentation results that are inconsistent with the semantics represented by the language, sometimes even diverging significantly. To this end, we propose a novel cross-modal conditioned Reconstruction for Language-guided Medical Image Segmentation (RecLMIS) to explicitly capture cross-modal interactions, which assumes that well-aligned medical visual features and medical notes can effectively reconstruct each other. We introduce conditioned interaction to adaptively predict patches and words of interest. Subsequently, they are utilized as conditioning factors for mutual reconstruction to align with regions described in the medical notes. Extensive experiments demonstrate the superiority of our RecLMIS, surpassing LViT by 3.74% mIoU on the publicly available MosMedData+ dataset and achieving an average increase of 1.89% mIoU for cross-domain tests on our QATA-CoV19 dataset. Simultaneously, we achieve a relative reduction of 20.2% in parameter count and a 55.5% decrease in computational load. The code will be available at https://github.com/ShashankHuang/RecLMIS.
@@
Figure 1: (a) Compared to previous work LViT [1], the proposed RecLMIS can focus on regions more matching the text. (b) We introduce cross-modal conditioned reconstruction to efficiently model visual-language alignment. (b) Compared to previous works (e.g., LViT [1]), RecLMIS does not require complex multi-level interactive alignment of features to achieve good results because it explicitly models the semantic connections between language and vision by automatically learning to reconstruct key features in text and images during training. (c) Comparison with state-of-the-art methods on MosMedData+ [2, 3, 1] on mIoU (y-axis), parameter count (size of the area), and FLOPs (x-axis).
@@ 1 Introduction
Medical image segmentation (MIS) is one of the most fundamental yet challenging tasks in medical image analysis. It plays a crucial role in numerous medical applications, such as disease identification [4], diagnosis [5], prognosis [6], and so on. However, owing to the typically expensive and challenging nature of obtaining expert-level segmentation annotations for medical images, there remains a scarcity of high-quality annotations in medical image segmentation. A feasible solution is to utilize medical notes to remedy the quality defects of medical image segmentation datasets [1]. In this paper, we follow [1] and focus on language-guided medical image segmentation, which is a challenging vision-language task.
@@
The one challenge is how to capture discriminative multi-modal representations for medical image segmentation. Although several recent approaches [1, 16] have attempted to incorporate textual features to improve medical image segmentation, they all rely on ambiguous attention mechanisms to embed text into visual features. Due to the highly similar foreground and background in medical images, this implicit modeling often leads to segmentation results that do not match the descriptions in medical notes. As shown in Fig. 1(a), we present the attention maps of the visual encoder. It is evident that the previous state-of-the-art method, LViT [1], struggled to accurately attend to the semantic content of “upper lower left lung” in the text, leading to incorrect prediction.
Another challenge lies in the relatively slow inference speed. Some methods achieve better integration of textual and visual semantics by introducing specific neural network layers. While this enhances performance, it also implies a greater number of parameters and computational load, consequently leading to slower inference speeds for the model. For instance, a dual-layer parallel U-shaped structure is introduced to integrate textual semantics in LViT [1].
To address this issue, we propose a novel cross-modal conditioned Reconstruction for Language-guided Medical Image Segmentation (RecLMIS) as shown in Fig. 1(b), which introduces conditioned reconstruction to explicitly capture cross-modal interactions of medical image and text.
Concretely, we assume that aligned text words and image regions reconstruct each other well. We first utilize the Conditional Interaction Module to score each token to adaptively capture patches of interest (POI) and words of interest (WOI).
Then, we treat language-guided medical image segmentation as a cross-modal reconstruction to explicitly investigate semantically relevant visual and textual concepts. POI/WOI are taken as probability distributions to mask images/texts randomly, and we consider WOI/POI as soft conditions to reconstruct images/texts.
Finally, we use the mean squared error (MSE) to measure the similarity between the reconstructed image/text and the original image/text. Simultaneously, we introduce a novel conditional contrastive learning loss to leverage POI and WOI to dominate cross-modal contrasts.
@@ 2 Related Work
With the advent of deep convolutional neural networks (CNN), the U-shape architecture was introduced for medical image segmentation in [17], which is improved by U-Net++ [18], UNet3+ [19], nnUnet [20], and so on. Recently, driven by the success of the Transformer and vision transformer, the researchers introduced two advanced combining CNN with transformer and pure transformer models. For example, Chen et al. [21] exploit both detailed high-resolution spatial semantic information from CNN features and the global context encoded by Transformers. Cao et al. [22] introduce a pure Transformer-based U-shaped Encoder-Decoder architecture with a skip connection. Furthermore, there is ongoing research [23, 24, 25, 26, 27] that delves into the development of specialized medical image Transformer structures, considering both encoder and decoder perspectives and attain notable performance. One of the pivotal factors contributing to the success of these models is that Transformer has an expansive global receptive field, endowing them with heightened flexibility and enhanced effect in addressing intricate patterns and relationships.
@@
Figure 2: Overview of the proposed cross-modal conditioned Reconstruction for Language-guided Medical Image Segmentation (RecLMIS). Given a pair of medical images and notes as prompts, we first exploit the visual encoder and text encoder to extract image and text features respectively. (d) The Conditioned Interaction (Sec. 3.2.1) module comprises a text-to-vision encoder and a vision-to-text encoder designed to align features from both vision and language inputs. (c, e) The conditioned vision-language reconstruction encompasses a Conditioned Vision Reconstruction module (CVR, Sec. 3.3.1) and a Conditioned Language Reconstruction module (CLR, Sec. 3.3.2), serving the purposes of medical image reconstruction and reconstruction of medical notes, respectively. (b) Finally, a lightweight mask predictor is employed to segment the designated region indicated by the text prompts.
@@ 2.2 Referring Image Segmentation
Compared to traditional single-modality visual segmentation tasks that operate under fixed category conditions, RIS must deal with the significantly more extensive lexicons and syntactic variations inherent in human natural languages. In conventional pipelines, there exist two primary processes: (1) extracting text features with language Transformers or recurrent neural networks, and extracting vision features with vanilla fully convolutional networks, and (2) building mask predictors with the fusion of these multi-modal features.
@@
In addition to that, many researchers have gradually focused on the first process, which has also achieved better performance, such as LAVT [36] and DMMI [11]. The former designs an early fusion scheme that effectively harnesses the Transformer encoder for modeling multi-modal context, and the latter reconstructs the masked entity phrase conditioned on the visual feature for facilitating the comprehensive understanding of the text expression. While visual-language segmentation models have made remarkable advancements in the realm of natural images, their direct application to medical image segmentation proves unsatisfactory due to the significant divergence between medical and natural images.
@@ 3 Method
Our RecLMIS adopts the encoder-decoder paradigm as shown in Fig. 2. The fundamental architecture for both encoder and decoder is based on the U-Net [17] structure with the PLAM [1] module. After the feature encoder, the Conditioned Interaction (Sec. 3.2.1) module with conditioned contrastive learning loss is employed for aligning visual and text features. Meanwhile, the Conditioned Language Reconstruction module (Sec. 3.3.2) with vision-to-text reconstruction loss, and the Conditioned Vision Reconstruction module (Sec. 3.3.1) with text-to-vision reconstruction loss are utilized to obtain a comprehensive cross-modal understanding of the textual expression and visual context. We elaborate on each component of our RecLMIS in detail in the following sections.
@@ 3.1 Feature Encoder
Given the input image $X$, we utilize the UNet [17] backbone with the PLAM [1] module as a visual encoder to extract visual features $V=\left\{v_{n}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ denote the number of channels, height, and width, respectively. The $N$ is equal to $H\times W$. For the medical note, we employ the frozen CLIP [7] text encoder to generate text features $E=\left\{e_{l}\right\}_{l=1}^{L}\in\mathbb{R}^{C\times L}$, where $L$ indicates the length of the medical note.
@@ 3.2 Conditioned Interaction Module
In human perception, when we recognize an image, our brain rapidly comprehends the accompanying textual information and integrates it with the visual content we observe, enabling us to pinpoint and localize the area of interest swiftly. However, in computer vision, the effective fusion of image and text features poses a challenging task. Several studies have already demonstrated the effectiveness of masking and reconstruction in enhancing representation learning [37, 38, 39, 40, 41]. These methods utilize various reconstruction techniques or visual generation methods to enhance the model’s representational capacity, thereby improving downstream task performance. We share a consistent motivation as these methods, but our method does not require destroying the structure of the original input and effectively introduces it for cross-modal interaction. Consequently, we propose the Conditioned Interaction module to jointly refine the visual feature $V$ and text features $E$ to achieve text-to-vision and vision-to-text alignment.
@@ 3.2.1 Hierarchical Structure
As illustrated in Fig.3, the Conditioned Interaction module is implemented in two parts: text-to-vision encoder and vision-to-text encoder. Compared to previous works (e.g., LViT [1]), RecLMIS does not require complex multi-level interactive alignment of features to achieve good results because it explicitly models the semantic connections between language and vision by automatically learning to reconstruct key features in text and images during training, more details in Sec. 3.3. Specifically, given the input text features $E$ and visual features $V$, we utilize cross-attention operation to perform cross-modal feature interaction. For the vision-to-text encoder, we regard $E$ as the query and $V$ as the key and value to obtain a new text feature $E^{\prime}$. Here, the visual feature $V$ is first flattened to $\mathbb{R}^{C\times N}$. Then the feature interaction is expressed as:
@@
Similarly, we consider $V$ as the query and $E$ as the key and value in the text-to-vision encoder to get a new visual feature $V^{\prime}$. Thus, Utilizing the final layer of coarse-grained image features within the backbone architecture enables us to capture the relationship between a single word and a patch of the image. More precisely, we can obtain the patches of interest $W_{poi}\!=\!\{w_{v}^{i}\}_{i=1}^{N}$ and the word of interest $W_{woi}\!=\!\{w_{e}^{j}\}_{j=1}^{L}$ by the following equation:
@@ 3.2.2 Conditioned Contrastive Learning
3.3 Conditioned Vision-Language Reconstruction
As Sec. 3.2 stated, an effective cross-modal semantic segmentation model must possess a deep comprehension of the association between text and images. Nonetheless, the feature alignment in existing models is often implicit, and it doesn’t explicitly demonstrate the model’s ability to comprehend the correlation between the given text and the corresponding image. Here, we assume that when provided with a pair consisting of text and an image featuring obscured regions, the model should have the capability to fill in the masked image features based on the textual information. Likewise, if presented with a pair of text with masked words and an image, the model should be able to impute the absent text features based on the image content. Hence, CVR (Sec. 3.3.1) and CLR (Sec. 3.3.2) are proposed to align text and image features explicitly.
3.3.1 Conditioned Vision Reconstruction
For visual features, we consider $W_{poi}$ obtained by the conditioned interaction module as the masking probability distribution. In other words, the visual feature $v_{i}$ with higher $w_{v}^{i}$ has a higher masking probability. Define the mask ratio as $\alpha_{v}\in[0,1]$, we then set the $m\!=\!\alpha_{v}\!\cdot\!N$ patch features to zero to generate the masked visual feature $V_{m}$. Specifically, we pinpoint the patches of most significance to the model by utilizing the $W_{poi}$ obtained from the Conditioned Interaction module through a probability distribution approach. We then modify the visual features by setting the features corresponding to these positions to zero, ultimately producing the masked visual features $V_{m}$. Ideally, a model with a comprehensive understanding of image and text semantics is able to reconstruct $V_{m}$ well based on complete text features $E$ and words of interest $W_{woi}$. Inspired by [44, 45], we introduce a conditioned reconstructor, denoted as $\Psi$, to assess the semantic alignment between the words of interest and the patches of interest. As shown in Fig. 2, our conditioned reconstructor $\Psi$ takes $W_{woi}$ as a condition, masked visual features $V_{m}$ as the query, and the textual feature $E$ as the key and value. We employ the dot-product attention mechanism to capture cross-modal interactions. The reconstructed vision features $\hat{V}=\{\hat{v}\}_{i=1}^{N}$ can be formulated as:

### v2

@@ Abstract
Recent developments underscore the potential of textual information in enhancing learning models for a deeper understanding of medical visual semantics. However, language-guided medical image segmentation still faces a challenging issue. Previous works employ implicit and ambiguous architectures to embed textual information. This leads to segmentation results that are inconsistent with the semantics represented by the language, sometimes even diverging significantly. To this end, we propose a novel cross-modal conditioned Reconstruction for Language-guided Medical Image Segmentation (RecLMIS) to explicitly capture cross-modal interactions, which assumes that well-aligned medical visual features and medical notes can effectively reconstruct each other. We introduce conditioned interaction to adaptively predict patches and words of interest. Subsequently, they are utilized as conditioning factors for mutual reconstruction to align with regions described in the medical notes. Extensive experiments demonstrate the superiority of our RecLMIS, surpassing LViT by 3.74% mIoU on the publicly available MosMedData+ dataset and achieving an average increase of 1.89% mIoU for cross-domain tests on our QATA-CoV19 dataset. Simultaneously, we achieve a relative reduction of 20.2% in parameter count and a 55.5% decrease in computational load. The code will be available at https://github.com/ShawnHuang497/RecLMIS.
@@ 1 Introduction
Medical image segmentation (MIS) is one of the most fundamental yet challenging tasks in medical image analysis.
Figure 1: (a) Existing methods (e.g., LViT[1]) face the issue of not fully and effectively adhering to the correct text prompts. For example, the text ’all left’ is not fully reflected in the attention maps in LViT [1]. Compared to that, the RecLMIS we proposed can focus on regions that match the text prompts properly. (b) Comparison with state-of-the-art methods on MosMedData+ [2, 3, 1] dataset on mIoU (y-axis), parameter count (size of the area), and FLOPs (x-axis). The RecLMIS we proposed is superior in both performance and FLOPs.
Figure 2: A comparative analysis of different model architectures for mask modeling and Language-guided Medical Image Segmentation (LMIS). The orange patch represents the masked region. The ’V-Encoder’, ’T-Encoder’, ’V-Decoder’, and ’T-Decoder’, represent vision encoder, text encoder, vision decoder, and text decoder, respectively. ’Attn’ and ’MSE’ stand for ’Attention’ and Mean Squared Error, respectively. (a) The MAE and BERT series mask modeling architectures involve masking the original image/sentence and subsequently using the vision decoder V-Decoder/T-Decoder for reconstruction. For achieving referring image segmentation, fine-tuning downstream task datasets is necessary. This approach encompasses methods such as [4, 5, 6]. (b) The parallel U-shape architecture (LViT [1], LAVT [7] integrates an additional parallel U-shape structure to the original segmentation model for processing and fusing text, aiming to align visual and textual features. (c) The dual-branch fusion architecture [8, 9] employs N layers of cross-attention in the text branch to align visual and textual features, minimizing the Mean Squared Error (MSE) between post-interaction and pre-interaction textual features. (d) Our cross-modal conditional reconstruction fusion architecture proposes using textual/visual features and conditions to reconstruct current visual/textual features during training, enabling the alignment of visual and textual features with just a single layer of cross-attention.
It plays a crucial role in numerous medical applications, such as disease identification [10], diagnosis [11], prognosis [12], and so on. However, owing to the typically expensive and challenging nature of obtaining expert-level segmentation annotations for medical images, there remains a scarcity of high-quality annotations in medical image segmentation. A feasible solution is to utilize medical notes to remedy the quality defects of medical image segmentation datasets [1]. In this paper, we follow [1] and focus on language-guided medical image segmentation, which is a challenging vision-language task.
@@
The one challenge is how to capture discriminative multi-modal representations for medical image segmentation. Although several recent approaches [1, 21] have attempted to incorporate textual features to improve medical image segmentation, they all rely on ambiguous attention mechanisms to embed text into visual features. Due to the highly similar foreground and background in medical images, this implicit modeling often leads to segmentation results that do not match the descriptions in medical notes. As shown in Fig. 1(a), we present the attention maps of the visual encoder. It is evident that the previous state-of-the-art method, LViT [1], struggled to accurately attend to the semantic content of “all left” in the text, leading to incorrect prediction.
Another challenge lies in the relatively slow inference speed. Some methods achieve better integration of textual and visual semantics by introducing specific neural network layers. While this enhances performance, it also implies a greater number of parameters and computational load, consequently leading to slower inference speeds for the model. For instance, a dual-layer parallel U-shape structure in Fig. 2(b) is introduced to integrate textual semantics in LViT [1]. Moreover, some Referring Image Segmentation (RIS) methods [8, 9] in the natural world, as shown in Fig. 2(c), also introduce multi-layer structures and complex interactions to fuse visual and linguistic features, which increase inference time and parameter count. This will be elaborated on in detail in Sec. 4.3.2).
To address this issue, we propose a novel cross-modal conditioned Reconstruction for Language-guided Medical Image Segmentation (RecLMIS) as shown in Fig. 2(d), which introduces conditioned reconstruction to explicitly capture cross-modal interactions of medical image and text. Concretely, we assume that aligned text words and image regions reconstruct each other well. We first utilize the Conditional Interaction Module to score each token to adaptively capture patches of interest ($W_{poi}$) and words of interest ($W_{woi}$). Then, we treat language-guided medical image segmentation as a cross-modal reconstruction to explicitly investigate semantically relevant visual and textual concepts. $W_{poi}$/$W_{woi}$ are taken as probability distributions to mask images/texts randomly, and we consider $W_{woi}$/$W_{poi}$ as soft conditions to reconstruct image/text features. Finally, we use the mean squared error (MSE) to measure the similarity between the reconstructed image/text and the original image/text features. Simultaneously, we introduce a novel conditional contrastive learning loss to leverage $W_{poi}$ and $W_{woi}$ to dominate cross-modal contrasts.
@@ 2 Related Work
With the advent of deep convolutional neural networks (CNN), the U-shape architecture was introduced for medical image segmentation in [22], which is improved by U-Net++ [23], UNet3+ [24], nnUnet [25], and so on. Recently, driven by the success of the Transformer and vision transformer, the researchers introduced two advanced combining CNN with transformer and pure transformer models. For example, Chen et al. [26] exploit both detailed high-resolution spatial semantic information from CNN features and the global context encoded by Transformers. Cao et al. [27] introduce a pure Transformer-based U-shape Encoder-Decoder architecture with a skip connection. Furthermore, there is ongoing research [28, 29, 30, 31, 32] that delves into the development of specialized medical image Transformer structures, considering both encoder and decoder perspectives and attain notable performance. One of the pivotal factors contributing to the success of these models is that Transformer has an expansive global receptive field, endowing them with heightened flexibility and enhanced effect in addressing intricate patterns and relationships.
@@
Figure 3: Overview of the proposed cross-modal conditioned Reconstruction for Language-guided Medical Image Segmentation (RecLMIS). Given a pair of medical images and notes as prompts, we first exploit the visual encoder and text encoder to extract image and text features respectively. The Conditioned Interaction (Sec. 3.2) module is designed to align features from both vision and language inputs. The conditioned vision-language reconstruction encompasses a Conditioned Vision Reconstruction module (CVR, Sec. 3.3) and a Conditioned Language Reconstruction module (CLR, Sec. 3.4), serving the purposes of medical image reconstruction and language feature reconstruction, respectively. Finally, a lightweight mask predictor is employed to segment the designated region indicated by the text prompts.
@@ 2.2 Referring Image Segmentation
Compared to traditional single-modality visual segmentation tasks that operate under fixed category conditions, Referring Image Segmentation (RIS) must deal with the significantly more extensive lexicons and syntactic variations inherent in human natural languages. In conventional pipelines, there exist two primary processes: (1) extracting text features with language Transformers or recurrent neural networks, and extracting vision features with vanilla fully convolutional networks, and (2) building mask predictors with the fusion of these multi-modal features.
@@
In addition to that, many researchers have gradually focused on the first process, which has also achieved better performance, such as LAVT [7, 41] and DMMI [9]. The former designs an early fusion scheme that effectively harnesses the Transformer encoder for modeling multi-modal context, and the latter reconstructs the masked entity phrase conditioned on the visual feature for facilitating the comprehensive understanding of the text expression. Moreover, many recent works [42, 8, 43, 44] have proposed complex feature fusion modules to achieve significant performance improvements. For example, Liu et al. [8] propose a multi-modal mutual attention module and Shah et al. [43] develop LQMFormer with the gaussian-enhanced multi-modal fusion module. While visual-language segmentation models have made remarkable advancements in the realm of natural images, their direct application to medical image segmentation proves unsatisfactory due to the significant divergence between medical and natural images.
@@ 3 Method
Our RecLMIS adopts the encoder-decoder paradigm as shown in Fig. 3. The fundamental architecture for both encoder and decoder is based on the U-Net [22] structure with the PLAM [1] module. After the feature encoder, the Conditioned Interaction (Sec. 3.2) module with conditioned contrastive learning loss is employed for aligning visual and text features. Meanwhile, the Conditioned Language Reconstruction module (Sec. 3.4) with vision-to-text reconstruction loss, and the Conditioned Vision Reconstruction module (Sec. 3.3) with text-to-vision reconstruction loss are utilized to obtain a comprehensive cross-modal understanding of the textual expression and visual context. We elaborate on each component of our RecLMIS in detail in the following sections.
@@ 3.1 Feature Encoder
Given an input image $X$, we utilize the UNet [22] backbone with the PLAM [1] module as a visual encoder to extract visual features $V=\left\{v_{n}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ denote the number of channels, height, and width, respectively. The $N$ is equal to $H\times W$. Given a medical note $T$, we use the BPE encoding algorithm to convert the sentence into a list of token IDs $T^{\prime}\in\mathbb{R}^{L}$, where $L$ is the number of tokens. Then, we use a word embedding layer to map the token ID list $T^{\prime}$ to high-dimensional vectors $T^{\prime\prime}\in\mathbb{R}^{L}$. Finally, we input $T^{\prime\prime}$ into a 12-layer standard transformer encoding block to extract the text feature representation vector $E=\left\{e_{l}\right\}_{l=1}^{L}\in\mathbb{R}^{C\times L}$.
@@ 3.2 Conditioned Interaction Module
In language-guided image segmentation tasks, the fusion of visual and textual features directly impacts the segmentation results. As shown in Fig. 1(a) in Sec. 1, if the visual and textual features are not well aligned, the text prompt will fail to influence the segmentation results, contradicting the initial purpose of using text to guide segmentation. Numerous studies have demonstrated that using cross-attention to fuse features is currently the optimal choice, and combining it with contrastive learning loss can better align visual and textual features [9, 4, 8]. Although these methods have achieved notable results, there is still room for optimization in model and loss design. As shown in Fig. 2(b) and (c), these methods employ multiple fusion layers, resulting in significant parameters and computational overhead during the inference stage. Additionally, the contrastive learning losses in some studies [9, 45, 46] are not sufficiently fine-grained. Therefore, we aim to compress the fusion layers to a single layer and extend the contrastive loss to a more fine-grained patch-word level to reduce the impact of non-essential text tokens and image patches.
@@ 3.2.1 Hierarchical Structure
As illustrated in Fig.4, the CI module is implemented in two parts: text-to-vision encoder and vision-to-text encoder. Compared to previous works (e.g., LViT [1]), RecLMIS does not require complex multi-level interactive alignment of features to achieve good results because it explicitly models the semantic connections between language and vision by automatically learning to reconstruct key features in text and images during training, more details in Sec. 3.3 and Sec. 3.4. Specifically, given the input text features $E$ and visual features $V$, we utilize cross-attention operation to perform cross-modal feature interaction. For the vision-to-text encoder, we regard $E$ as the query and $V$ as the key and value to obtain a new text feature $E^{\prime}$. Here, the visual feature $V$ is first flattened to $\mathbb{R}^{C\times N}$. Then the feature interaction is expressed as:
@@
Figure 5: The structure of (a) Conditioned Vision Reconstruction module and (b) Conditioned Language Reconstruction.
Similarly, we consider $V$ as the query and $E$ as the key and value in the text-to-vision encoder to get a new visual feature $V^{\prime}$. Thus, Utilizing the final layer of coarse-grained image features within the backbone architecture enables us to capture the relationship between a single word and a patch of the image. More precisely, we can obtain the patches of interest $W_{poi}\!=\!\{w_{v}^{i}\}_{i=1}^{N}$ with shape $1\times N$ and the word of interest $W_{woi}\!=\!\{w_{e}^{j}\}_{j=1}^{L}$ with shape $1\times L$ by the following equation:
@@ 3.2.2 Conditioned Contrastive Learning
$\mathcal{S}_{v,e}=\sum_{j=1}^{L}w_{e}^{j}a_{ij},$ (4)
where $w_{e}^{j}$ denotes the attention weight of the $j$-th word. Although this method finely aligns image patches to each text token, it does not consider the relationship between each text token and every image patch in a fine-grained manner. Additionally, noisy regions in the text or image might interfere with the similarity calculation.
Inspired by fine-grained semantic alignment [47], we maintain the sequence structure of both the image and text. We introduce $w_{e}^{j}$ to calculate the contribution of each image patch to the similarity calculation. Additionally, to mitigate the impact of noise, we select only the token/patch with the highest similarity to participate in the calculation. The similarity between an image-text pair in our $\mathcal{L}_{CCL}$ calculation formula can be expressed as follows:
@@
3.3 Conditioned Vision Reconstruction
Several studies have demonstrated the effectiveness of masking and reconstruction in enhancing representation learning [49, 6, 50, 51, 52, 4], utilizing various reconstruction techniques or visual generation methods to improve model representational capacity and downstream task performance. These methods all fall under the category of Fig. 2(a), which involves masking and reconstructing the image itself. While such approaches can enhance the model’s text-to-vision representation capability, reconstructing the original image includes unimportant regions, thus adding unnecessary computational complexity. We aim to introduce an auxiliary module that, without altering the original model structure, utilizes text features $E$ to reconstruct masked key areas of the visual features $V_{m}$, thereby achieving text-to-vision alignment.
Method Backbone Text Parameters (M) FLOPs (G) QaTa-COV19 MosMedData+
Dice(%$\uparrow$ mIoU(%$\uparrow$ Dice(%$\uparrow$ mIoU(%$\uparrow$
U-Net [22] CNN 14.8 50.3 79.02 69.46 64.60 50.73
U-Net++ [23] CNN 74.5 94.6 79.62 70.25 71.75 58.39
nnUNet [25] CNN 19.1 412.7 80.42 70.81 72.59 60.36
TransUNet [26] Hybrid 105.0 56.7 78.63 69.13 71.24 58.44
Swin-Unet [27] Transformer 82.3 67.3 78.07 68.34 63.29 50.19
ConVIRT [53] CNN 35.2 44.6 79.72 70.58 72.06 59.73
TGANet [21] CNN 19.8 41.9 79.87 70.75 71.81 59.28
CLIP [13] Transformer 87.0 105.3 79.81 70.66 71.97 59.64
GLoRIA [46] CNN 45.6 60.8 79.94 70.68 72.42 60.18
ViLT [54] Hybrid 87.4 55.9 79.63 70.12 72.36 60.15
LAVT [7] Transformer 118.6 83.8 79.28 69.89 73.29 60.41
LViT [1] Hybrid 29.7 54.1 83.66 75.11 74.57 61.33
SLViT [55] Hybrid 131.5 51.1 79.25 68.87 72.57 60.78
DMMI [9] Transformer 114.6 63.3 84.13 75.66 75.01 61.83
RefSegformer [41] Transformer 195.0 103.6 84.09 75.48 74.98 61.70
RecLMIS CNN 23.7 24.1 85.22 77.00 77.48 65.07
As shown in Fig. 5(a), we consider $W_{poi}$ obtained by the CI module as the masking probability distribution. In other words, the visual feature $v_{i}$ with higher $w_{v}^{i}$ has a higher masking probability. Define the mask ratio as $\alpha_{v}\in[0,1]$, we then set the $m\!=\!\alpha_{v}\!\cdot\!N$ patch features to zero to generate the masked visual feature $V_{m}$. Specifically, we pinpoint the patches of most significance to the model by utilizing the $W_{poi}$ obtained from the CI module through a probability distribution approach. We then modify the visual features by setting the features corresponding to these positions to zero, ultimately producing the masked visual features $V_{m}$. Ideally, a model with a comprehensive understanding of image and text semantics is able to reconstruct $V_{m}$ well based on complete text features $E$ and words of interest $W_{woi}$. Inspired by [56, 57], we introduce a conditioned reconstructor, denoted as $\Psi$, to assess the semantic alignment between the words of interest and the patches of interest. As shown in Fig. 3, our conditioned reconstructor $\Psi$ takes $W_{woi}$ as a condition, masked visual features $V_{m}$ as the query, and the textual feature $E$ as the key and value. We employ the dot-product attention mechanism to capture cross-modal interactions. The reconstructed vision features $\hat{V}=\{\hat{v}\}_{i=1}^{N}$ can be formulated as: